### v1

# Boosting Open-Domain Continual Learning via Leveraging Intra-domain Category-aware Prototype

Abstract
The task of open-domain continual learning in Vision-Language Model (VLM) has become both a practical and yet challenging problem. In the Open-Domain Continual Learning (ODCL) task, VLMs are required to incrementally learn both seen and unseen categories from diverse datasets, enhancing the recognition of new categories while maintaining the ability to recognize previously seen categories. As Contrastive LanguageImage Pre-training (CLIP) is a VLM trained on a large number of image-text pairs, it possesses strong zero-shot extraction capabilities for features across different domains. However, existing methods addressing the ODCL task often overlook the inherent domain gaps. In this work, we extract domain prior information using a training-free method to enhance the performance of CLIP, named DPeCLIP. Specifically, we leverage the strong domain information extraction capabilities of the CLIP model to achieve training-free Domain Prior Classifier across different domains. Additionally, we introduce the Domain Prior Prompt method, which incorporates domain prior information into the model to bridge the significant gaps between domains. Extensive experiments conducted on 11 datasets demonstrate that our approach outperforms the current state-of-the-art methods, validating its effectiveness.
Figure 1: (a) Comparison of two approaches for solving the ODCL-CIL task: one approach involves using all seen categories for the classification of test images, while the other method employs a Task-ID discriminator to determine the domain corresponding to the test image, and then selects the categories of the corresponding domain for classification. (b) Comparison of Task-ID classification accuracy between MoE-Adapters and our method. (c) Comparison of the performance of our method and other methods in ODCL-CIL task on the Caltech101 and CIFAR100 datasets. The ODCL task requires evaluating both seen and unseen datasets. The black dashed line on the left indicates that the test dataset has not been trained on, and it is assessed through the model’s zero-shot capability.
Introduction
Continual learning (CL) (dhar2019learning; rebuffi2017icarl) aims to enable models to continually acquire new knowledge, such as learning to recognize new object categories, while addressing the catastrophic forgetting of old knowledge. In recent years, CL has gained significant attention and has been applied in various fields, including autonomous driving and medical diagnosis. However, traditional continual learning methods are largely limited to single-domain data, failing to effectively address the complexities of real-world scenarios. In practice, data often comes from multiple domains, and there may be significant differences between these domains.
To address this issue, open-domain continual learning (ODCL)  (zheng2023preventing; li2024coleclip) has been explored by many recent researchers (zheng2023preventing; yu2024boosting; li2024coleclip). The ODCL task utilizes Vision-Language Model (VLM) (jia2021scaling; li2022blip; radford2021learning) to linearly learn data from different domains, acquiring new domain knowledge while preventing the forgetting of old domain knowledge. Due to the zero-shot classification capabilities of VLM, ODCL is tested on both seen and unseen domains, necessitating the preservation of VLM’s original zero-shot classification abilities. Similar to traditional continual learning, the ODCL task includes both class-incremental (CIL) and task-incremental (TIL) settings. Compared to ODCL-TIL settings, ODCL-CIL does not have access to the Task-ID of the current test image during inference.
ODCL task faces two significant challenges: (1) Two types of Forgetting: Unlike traditional CL task, ODCL task faces not only the forgetting of old knowledge but also the forgetting of zeroshot capabilities. (2) Influence of Category-relatedness Between Domains: For example, although ”wheelchair” in Caltech101 and ”chair” in CIFAR100 belong to entirely different categories, their semantic similarity can cause confusion in the ODCL-CIL task, leading to a decrease in dataset performance.
Currently, there are two primary technical approaches to addressing ODCL task, as illustrated in Figure 1(a). For the first approach, during the inference stage, the model utilizes the embeddings of all seen categories to classify the test image regardless of which domain the test image comes from. ZSCL (zheng2023preventing) leverages a large reference dataset combined with the distillation of original CLIP knowledge. Despite the distillation, fully fine-tuning CLIP results in significant forgetting. ColeCLIP (li2024coleclip) tries to mitigate the interference of similar categories across different domains by maintaining a large vocabulary and employing a finely designed task prompt. But, since ColeCLIP utilizes the same text embedding for the same category across different datasets (domains), although momentum is used to slow down changes in text embeddings during this process, the lack of updates to the old task prompt leads to inconsistency between the text embedding and the old task prompt. As shown in Figure 1(c), with the progression of training stages, ColeCLIP exhibits significant catastrophic forgetting issues. Besides, its zeroshot performance also declines. These methods try to alleviate the forgetting problem, but since the model cannot access data from multiple domains simultaneously during training, and similar categories from different domains can cause confusion during testing, addressing the category-relatedness problem becomes very challenging.
The second approach involves first determining the Task-ID of the test image using a Task-ID discriminator, and then classifying it using the category set corresponding to the identified Task-ID. For example, MoE-Adapters (yu2024boosting) first propose a combination of Auto-Encoder (AE) and AlexNet (krizhevsky2012imagenet) to determine the Task-ID followed by exploring the performance of CLIP on ODCL task by adopting a Mixture of Experts (MoE) (jacobs1991adaptive) structure. Despite the expert freezing mechanism, the continual updating of experts still results in knowledge forgetting. As shown in Figure 1(b) and (c), the effectiveness of the Task-ID discriminator directly affects the model’s final performance. Due to the limited performance of the Task-ID discriminator, the overall performance of the model is affected.
Compared to the first approach, the second approach offers a better solution to addressing the issue of the category-relatedness, namely that a more robust Task-ID discriminator can more effectively resolve this problem. Besides, we find that even with a finely designed domain prompt, updating it using cross-domain information still fails to alleviate the forgetting issue (li2024coleclip), since the task prompt can still be influenced by the inconsistency between text embeddings and old task prompt. Therefore, if the following two factors can be considered, it will greatly solve the category-relatedness problem and mitigate the forgetting of old domains: (1) correctly identifying the Task-ID of a test image, and using only the category set corresponding to the Task-ID, and (2) preserving the knowledge only related to each domain. Additionally, the model should retain the original knowledge of CLIP to maintain a good zeroshot capability.
Based on the above considerations, we propose a simple yet effective solution: leveraging intra-domain category-aware prototypes for ODCL in CLIP (radford2021learning) framework, dubbed as Domain Prototype enhanced CLIP (DPeCLIP). To maintain the original zeroshot capability of CLIP, we propose to average the original outputs of CLIP’s image and text modalities belonging to the same category within a domain as prototypes. We propose to (1) distinguish the Task-ID of a test image, and (2) preserve the knowledge only related to each domain, where prototypes is the key to bridging them. Firstly, we propose a robust, training-free Task-ID discriminator, by utilizing prototypes as classifiers for identifying Task-IDs. Compared to MoE-Adapters, our method significantly improves Task-ID accuracy without introducing additional training parameters. Secondly, to maintain the knowledge corresponding to each domain during training, we incorporate intra-domain category-aware prototypes as domain prior prompts into the training process. For the text branch, we propose a text self-attention module which encodes the relationships of categories within a domain as prompts into the text branch. For the image branch, unlike the text branch, we introduce an image cross-attention module which uses instance-level image embeddings to query category-level prototypes. This ensures that the instance prompt carries information about the relationships between various categories, enabling the learned prompt to encapsulate both domain and instance information.
Overall, our contributions can be summarized below:
  • We propose a robust, training-free Task-ID discriminator method that utilizes domain-specific, category-aware prototypes as classifiers to effectively distinguish test images from the original domain. This method not only addresses the issue of strong category-relatedness across different domains, but also significantly improves the model’s performance on ODCL task.
  • We incorporate category-domain class-aware prototypes as domain prior prompts into the training process to maintain the knowledge corresponding to each domain. This method not only enhances the model’s recognition capability for new categories, but also maintains good zero-shot classification ability.
  • Through comprehensive experiments on 11 datasets, we demonstrate that our method achieves state-of-the-art (SOTA) performance in both ODCL-CIL and ODCL-TIL settings, compared to previous best methods.
Related Work
Continual Learning
To address the issue of catastrophic forgetting, continual learning can be broadly categorized into three primary approaches: replay-based methods (bang2021rainbow; chaudhry2018riemannian; rebuffi2017icarl; lopez2017gradient; prabhu2020gdumb; shin2017continual), regularization-based methods (kirkpatrick2017overcoming; zenke2017continual; li2017learning; dhar2019learning; douillard2020podnet; hou2019learning), and parameter expansion-based methods (gao2023unified; smith2023coda; wang2022foster; yan2021dynamically; zhou2022model; wang2022dualprompt; wang2022learning; wang2022s). Replay-based methods maintain a buffer of previously encountered data and leverage this replay buffer during the learning of new data to retain the original knowledge. Regularization-based methods aim to preserve prior knowledge by constraining the direction of model updates, with knowledge distillation (hinton2015distilling) being a prevalent technique. Parameter expansion-based methods mitigate forgetting by expanding the model’s architecture to accommodate new tasks. Recently, several parameter-efficient fine-tuning (PEFT) based on parameter expansion methods have emerged, such as Low-Rank Adaptation (LoRA) (hu2021lora), prompt-based methods (lester2021power), and adapter modules (houlsby2019parameter). L2P (wang2022learning) addresses the issue of preserving old knowledge by maintaining a large prompt pool. DualPrompt (wang2022dualprompt) further mitigates catastrophic forgetting by employing both general and specific prompts. CODA-Prompt (smith2023coda) extends this approach by increasing the scale of prompts and utilizing more prompts simultaneously through a weighting mechanism. LAE (gao2023unified) explores the use of three PEFT methods by implementing both online and offline update strategies.
Continual Learning of CLIP
Currently, numerous studies focus on the performance of CLIP in continual learning, which can be divided into two categories: traditional continual learning (ding2022don; zhou2023learning; wang2022s) and ODCL tasks (zheng2023preventing; yu2024boosting; li2024coleclip).
For traditional continual learning with CLIP, VR-LwF (ding2022don) retains old knowledge through a simple vocabulary distillation method, Proof (zhou2023learning) employs a cross-modal fusion module to fully utilize knowledge and task-specific mapping layers to alleviate forgetting, and S-Prompt (wang2022s) uses k-means to save the center of each domain as a Task-ID discriminator, selecting the corresponding domain’s prompt for test images.
ZSCL (zheng2023preventing), as the first method proposed for open-domain tasks, balances newly learned knowledge and old knowledge through feature-level distillation with an external dataset. MoE-Adapters (yu2024boosting) improve performance in ODCL task by introducing an external dataset to train corresponding autoencoders for Task-ID identification and using a MoE structure. CoLeCLIP (li2024coleclip) extends open-domain tasks, by maintaining a vocabulary and designing skillfully task prompts to store task-specific knowledge, but this approach inevitably leads to performance degradation in ODCL-CIL task. In comparison, our method achieves significant performance improvement by determining the Task-ID through a Domian Prior Classifier approach.
Figure 2: Domain Prototype enhanced CLIP (DPeCLIP) framework.The three stages are presented as follows: (a) Prototype Calculation Stage, where category-aware prototypes are extracted using the original CLIP model. (b) Training Stage, where we utilize two sub-modules, Text Self-Attention (TSA) and Image Cross-Attention (ICA), to provide Domain Prior Prompts for different categories and images, conditioned on the domain prototypes as input. Additionally, we extract domain-level shared information using general prompts. (c) Inference Stage, where we use the Domain Prototypes to determine the Task-ID for test images and employ the corresponding Domain Components and category sets associated with the Task-ID for classifying the test images.
Methodology
Preliminaries
CLIP
CLIP is a model endowed with zero-shot classification capabilities, comprising an image encoder $\mathcal{I}(\cdot)$ and a text encoder $\mathcal{T}(\cdot)$. The image encoder can be configured as either ViT (dosovitskiy2020image) or ResNet (he2016deep), whereas the text encoder is based on the Transformer architecture (vaswani2017attention). These encoders are trained using contrastive loss on a vast amount of paired image-text samples, facilitating a high degree of alignment between image and text features, which imparts robust zero-shot classification abilities to the model.
To employ the CLIP model for classification tasks, an image $x$ is first processed through the image encoder to extract image features. The target classes $\{c^{1},c^{2},\cdots,c^{J}\}$ for discrimination are converted into corresponding textual descriptions $\{\hat{c}^{1},\hat{c}^{2},\cdots,\hat{c}^{J}\}$ using a template such as “a photo of a”, where $J$ represents the total number of categories. These textual descriptions are then fed into CLIP text encoder to generate text features. The cosine similarity between the text features and the image features is computed, and the candidate category with the highest similarity is selected as the final classification result.
ODCL
For the ODCL tasks, the model is required to sequentially learn from $N$ distinct datasets (domains). The primary objective of this task is to enhance the recognition of new categories while maintaining the ability to recognize previously learned categories. Given $N$ datasets, $\{\mathcal{D}_{1},\mathcal{D}_{2},\cdots,\mathcal{D}_{N}\}$, as the model learns these datasets linearly, after learning the $n$th dataset $\mathcal{D}_{n}$, it exhibits good classification capability on the previously learned datasets $\{\mathcal{D}_{1},\mathcal{D}_{2},\cdots,\mathcal{D}_{n}\}$. At the same time, the model aims to preserve its zero-shot classification ability for the unlearned datasets $\{\mathcal{D}_{n+1},\mathcal{D}_{n+2},\cdots,\mathcal{D}_{N}\}$.
To thoroughly evaluate the model’s performance on ODCL, there are two settings for inference: task-incremental learning and class-incremental learning, denoted by ODCL-TIL and ODCL-CIL, respectively. In the ODCL-TIL task, the Task-ID is known during inference, allowing us to utilize the category set corresponding to the test image. In contrast, in the ODCL-CIL task, the Task-ID is not accessible during inference, necessitating the use of all seen categories to classify the test image. Clearly, for ODCL-CIL task, the inability to obtain the task ID makes it a more challenging task compared to ODCL-TIL.
Overview of DPeCLIP
Figure 2 illustrates the overall architecture of our method. Our method consists of three parts: (a) Prototypes Calculation Stage, (b) Training Stage, and (c) Inference Stage. In the Prototypes Calculation Stage, we utilize the CLIP image encoder to extract the image features of training images belonging to $j$th category within $n$th domain and average them to obtain the mean image feature $\bar{I}_{n}^{j}$. Likewise, we use the CLIP text encoder to extract the text features, combined with CLIP templates, such as “a photo of a”. By averaging these text features, we obtain the average text feature $\bar{T}_{n}^{j}$. Finally, by performing element-wise summation of $\bar{I}_{n}^{j}$ and $\bar{T}_{n}^{j}$, we derive one intra-domain category-aware Prototype $P_{n}^{j}$.
During the training stage, we propose Domain Prior Prompts through two sub-modules: Text Self-Attention (TSA) and Image Cross-Attention (ICA), conditioned on the prototypes as input. In addition, we also learn the shared information for each domain through the General Prompts. In this process, both domain prior prompt and the general prompt are updated using cross-entropy loss.
During the inference stage, we first utilize the domain prototypes to determine the Task-ID. Then, based on the identified Task-ID, we select the corresponding TSA, ICA modules, General Prompts, and the category set associated with the Task-ID from the domain prior prompt pool, which contains all trained components, to classify the test image.
Prototype Calculation Stage
To obtain intra-domain category-aware prototypes for ODCL, we extract prototypes from the training set using the original CLIP model, as shown in Figure 2(a). Specifically, for the $j$th category in $\mathcal{D}_{n}$, containing $M$ training images, the image prototype $\bar{I}_{n}^{j}$ can be calculated via:
$\bar{I}_{n}^{j}=\frac{\sum_{m=1}^{M}\mathcal{I}(x_{n}^{j,m})}{M},$ (1)
where $x_{n}^{j,m}$ means the $m$th image belonging to the $j$th category of $\mathcal{D}_{n}$. We adopt different templates provided by CLIP to generate various textual descriptions for a category name. For the $j$th category in the $n$th domain $c_{n}^{j}$, the textual description generated by the $z$th template is denoted as $\hat{c}_{n}^{j,z}$. Then the text prototype $\bar{T}_{n}^{j}$ is calculated by:
$\bar{T}_{n}^{j}=\frac{\sum_{z=1}^{Z}\mathcal{T}(\hat{c}_{n}^{j,z})}{Z},$ (2)
where $Z$ represents the total number of templates.
Subsequently, we obtain the intra-domain category-aware prototypes $P_{n}^{j}$ by performing element-wise summation of $\bar{I}_{n}^{j}$ and $\bar{T}_{n}^{j}$.
Training Stage
Text Branch
For the text encoder $\mathcal{T}(\cdot)$, we provide the relationships between different categories within the domain through the domain prior prompt, while learning the general information of the domain through the general prompt, thereby enhancing the model’s classification performance, as shown in Figure 2(b).
We denote the category name of the $j$th category in $\mathcal{D}_{n}$ as $c_{n}^{j}$. $c_{n}^{j}$ is then processed through CLIP word embedding layer to obtain the text tokens $[t_{1}^{(1)},t_{2}^{(1)},\cdots,t_{l}^{(1)}]$, subscript $l$ represents the length of text tokens and the superscript (1) represents the first layer. In the first layer $\mathcal{T}^{(1)}(\cdot)$ of the text encoder, we use a general prompt $v_{g}^{(1)}$, where $v_{g}^{(1)}\in\mathbb{R}^{1\times d^{\mathcal{T}}}$ and $d^{\mathcal{T}}$ is the dimension of the text token. This prompt is concatenated with the text tokens of different categories. In this process, only $v_{g}^{(1)}$ is learnable. Then the output of the first text encoder layer is:
$[t_{1}^{(2)},t_{2}^{(2)},\cdots,t_{l}^{(2)},v_{g}^{(2)}]=\mathcal{T}^{(1)}([t_{1}^{(1)},t_{2}^{(1)},\cdots,t_{l}^{(1)},v_{g}^{(1)}]).$ (3)
Moreover, we propose a category-aware domain prior prompt for the text encoder. Specifically, conditioned on the prototype $P_{n}=[{P_{n}^{1}}^{\top},{P_{n}^{2}}^{\top},\cdots,{P_{n}^{S}}^{\top}]^{\top}$ as input, where $P_{n}\in{\mathbb{R}^{S\times d^{\mathcal{T}}}}$. We propose a Text Self-Attention (TSA) module to generate the domain prior prompt $v_{p}$, $v_{p}\in{\mathbb{R}^{S\times d^{\mathcal{T}}}}$. The TSA module is depicted in Figure 2 (with Add and Norm omitted). We use the domain prototype $P_{n}$ as the input for ${Q}$, ${K}$, and ${V}$ in TSA, thereby obtaining $v_{p}$.
$Q=W_{Q}\cdot P_{n},K=W_{K}\cdot P_{n},V=W_{V}\cdot P_{n},$ (4)
$\mathtt{Attention}(Q,K,V)=\mathtt{softmax}(\frac{QK^{\top}}{\sqrt{d^{\mathcal{T}}}})\cdot V,$ (5)
$v_{p}=\mathtt{FFD}(\mathtt{Attention}(Q,K,V)),$ (6)
where $W_{Q}$, $W_{K}$ and $W_{V}\in\mathbb{R}^{d^{\mathcal{T}}\times d^{\mathcal{T}}}$ are learnable projection matrices, and $\mathtt{FFD}$ represents Feed Forward layer.
These domain prior prompts $v_{p}$ then replace the general prompt in the $h$-th layer of the text encoder. Thus, the output of the $h$th text encoder layer is:
$\begin{split}[t_{1}^{(h+1)},t_{2}^{(h+1)},\cdots,t_{l}^{(h+1)},v_{p}^{(h+1)}]\\ =\mathcal{T}^{(h)}([t_{1}^{(h)},t_{2}^{(h)},\cdots,t_{l}^{(h)},v_{p}]).\end{split}$ (7)
Image Branch
As shown in Figure 2(b), similar to the text branch, we use the general prompt to learn the general information of the domain for image encoder $\mathcal{I}(\cdot)$. In addition, we utilize the ICA module to learn the instance-level domain prior prompt, thereby providing both domain and instance information simultaneously. Specifically, the image passes through the patch embedding layer to obtain image tokens $[cls^{(1)},i_{1}^{(1)},\cdots,i_{l^{\prime}}^{(1)}]$, where $cls^{(1)}$ is class token in the first layer and subscript $l^{\prime}$ is the number of image tokens. In the first $h-1$ layers of the image encoder, we concatenate a general prompt $e_{g}$, where $e_{g}\in{\mathbb{R}^{1\times d^{\mathcal{I}}}}$ with the image tokens at each layer, and $d^{\mathcal{I}}$ means the dimension of image token. For example, for the $h-1$ layer $\mathcal{I}^{(h-1)}(\cdot)$ of image encoder, the general prompt $e_{g}^{(h-1)}$ is used by following formula.
$\begin{split}&[cls^{(h)},t_{1}^{(h)},\cdots,t_{l^{\prime}}^{(h)},e_{g}^{(h)}]=\\ &\mathcal{I}^{(h-1)}([cls^{(h-1)},t_{2}^{(h-1)},\cdots,t_{l^{\prime}}^{(h-1)},e_{g}^{(h-1)}])\end{split}$ (8)
In this process, all $e_{g}^{(1)},e_{g}^{(2)},...,e_{g}^{(h-1)}$ are learnable.
In addition to the general prompt, we provide instance-level domain prior prompt for the image encoder. Using the prototype $P_{n}$ as input, we propose an Image Cross-Attention (ICA) module to dynamically generate instance-level prompt. The class token ${cls}^{(h)}$ from the $h$th layer of the image encoder serves as the $Q$, while the domain prototype serves as $K$ and $V$, resulting in the domain prior prompt $e_{p}$, $e_{p}\in{\mathbb{R}^{1\times d^{\mathcal{I}}}}$.
$Q^{\prime}=W_{Q}^{\prime}\cdot{cls}^{(h)},K^{\prime}=W_{K}^{\prime}\cdot P_{n},V^{\prime}=W_{V}^{\prime}\cdot P_{n},$ (9)
where $W_{Q}^{\prime}\in\mathbb{R}^{d^{\mathcal{I}}\times d^{\mathcal{I}}}$, $W_{K}^{\prime}$ and $W_{V}^{\prime}\in\mathbb{R}^{d^{\mathcal{T}}\times d^{\mathcal{I}}}$ are learnable projection matrices in the image branch.
After training each domain, we save the general prompts, TSA, and ICA modules as the domain component into the domain prior prompt pool for inference.
Inference Stage
For test images from previously seen domains in the ODCL-CIL setting, we first employ the prototype Task-ID discriminator to determine the Task-ID of the test image, as shown in Figure 2(c). Specifically, for a test image $x$, we use the original CLIP image encoder to extract the corresponding image feature $I$.
$I=\mathcal{I}(x).$ (10)
We then calculate the cosine similarity between the image feature $I$ and the learned domain prototypes of different domains. For each domain, we select the highest similarity score, and then compare the similarities scores across different domains to obtain the Task-ID. Subsequently, we select the corresponding domain component from the Domain prior prompt pool. Classification of the test image is then performed within the category set associated with the identified Task-ID domain and domain component.
For test images from unseen domains, we adopt an approach similar to ColeCLIP. We utilize CLIP to perform zero-shot classification on these images from unseen domains. However, if the unseen dataset contains some categories that have been previously seen, we will query the previously seen domains in which the corresponding categories appear. We will then utilize the components associated with those domains to compute the cosine similarity for the respective categories. The category with the highest similarity will be selected as the corresponding result. Finally, we will compare this with the cosine similarity of CLIP for unseen categories, choosing the largest as the final classification result.
Method Aircraft Caltech101 CIFAR100 DTD EuroSAT Flowers Food MNIST OxfordPet Cars SUN397 Mean
Last CODA-Prompt 12.81 65.38 63.49 53.83 56.91 56.64 83.31 83.99 76.83 53.65 75.13 62.00
LAE 13.65 83.06 66.88 59.20 36.61 48.95 84.08 94.22 75.44 42.69 78.69 62.14
CLIP 24.39 63.65 40.99 39.26 52.98 70.04 88.40 39.56 88.85 64.52 63.28 57.81
ZSCL 42.54 64.40 67.24 54.84 89.70 90.37 91.73 95.76 93.38 85.20 78.30 77.59
MoE-Adapters 34.11 47.58 80.88 75.48 0.00 93.02 70.78 99.35 86.40 79.80 68.89 66.93
CoLeCLIP 48.09 73.10 65.22 69.63 83.98 96.21 90.88 94.56 93.54 82.61 79.34 79.74
DPeCLIP 49.86 84.21 83.23 71.06 97.01 95.77 92.02 99.40 93.89 84.47 80.17 84.64
Forgetting CODA-Prompt 20.18 73.61 73.43 62.75 84.66 69.79 87.42 83.00 80.96 65.57 75.13 70.59
LAE 17.07 85.58 75.64 67.70 64.90 70.45 87.63 96.64 86.54 62.49 78.69 72.12
CLIP 24.42 66.26 48.92 41.03 53.09 70.04 88.43 39.56 88.93 64.53 63.28 58.95
ZSCL 46.32 66.72 75.87 60.61 94.85 91.75 91.83 94.30 94.14 85.48 78.30 80.02
MoE-Adapters 37.17 63.06 82.26 76.18 0.00 93.60 71.03 99.42 86.97 80.15 68.89 68.97
CoLeCLIP 48.24 76.72 73.15 73.51 89.74 96.21 91.02 94.56 93.76 82.64 79.34 81.72
DPeCLIP 49.86 84.97 84.63 72.96 97.36 95.77 92.03 99.40 93.93 84.48 80.17 85.05
Avg CLIP 24.42 68.21 52.28 42.14 53.66 70.27 88.58 52.20 89.07 64.55 65.25 60.97
ZSCL 46.32 68.34 74.34 56.29 79.10 81.42 89.52 73.96 89.01 64.35 67.49 71.83
MoE-Adapters 37.17 65.34 79.45 67.59 19.66 83.13 80.53 73.98 88.48 67.51 65.34 66.19
CoLeCLIP 48.24 77.76 71.69 65.65 76.78 83.76 89.64 72.22 90.33 67.97 66.40 73.68
DPeCLIP 49.86 85.26 81.45 65.25 81.58 84.31 89.93 73.97 90.37 68.30 66.17 76.05
Table 1: $Last$, $Forgetting$, and $Avg$ results(%) in ODCL-CIL. CODA-Prompt and LAE are not applicable to the Avg metric.
Method Aircraft Caltech101 CIFAR100 DTD EuroSAT Flowers Food MNIST OxfordPet Cars SUN397 Mean
Last CODA-Prompt 43.59 92.97 74.47 75.21 92.57 93.07 90.25 98.61 92.75 83.80 78.75 83.28
LAE 49.14 93.84 80.53 75.43 90.63 87.07 88.94 98.64 92.94 81.36 80.08 83.51
CLIP 24.42 87.79 67.36 45.11 54.65 70.53 88.70 59.43 89.13 64.56 65.45 65.19
ZSCL 42.60 92.40 81.37 70.64 94.63 90.58 91.76 98.64 93.70 85.21 79.69 83.75
MoE-Adapters 34.59 94.70 82.69 76.91 97.65 94.76 91.86 99.35 94.74 80.86 80.54 84.42
CoLeCLIP 48.66 94.93 78.78 78.35 88.85 96.29 91.10 97.63 94.39 82.73 80.15 84.71
DPeCLIP 49.86 95.62 85.79 78.61 98.38 95.77 92.11 99.40 93.97 84.49 81.69 86.89
Frogetting CODA-Prompt 47.49 92.84 78.23 76.40 94.51 95.23 90.20 98.55 93.12 83.99 78.75 84.48
LAE 53.69 94.06 80.50 76.72 93.54 90.34 89.46 98.97 93.42 84.38 80.08 85.02
CLIP 24.42 87.79 67.36 45.11 54.65 70.53 88.70 59.43 89.13 64.56 65.45 65.19
ZSCL 46.35 93.45 82.67 71.56 96.15 91.98 91.84 98.78 94.36 85.49 79.69 84.76
MoE-Adapters 37.38 94.41 83.56 77.17 97.79 95.32 91.88 99.42 94.57 81.01 80.54 84.82
CoLeCLIP 48.66 94.96 79.16 78.35 93.25 96.29 91.10 97.63 94.39 82.73 80.15 85.15
DPeCLIP 49.86 95.62 85.79 78.61 98.38 95.77 92.11 99.40 93.97 84.49 81.69 86.89
Avg CLIP 24.42 87.79 67.36 45.11 54.65 70.53 88.70 59.43 89.13 64.56 65.45 65.19
ZSCL 46.35 92.64 79.91 64.26 79.93 81.54 89.53 75.59 89.07 64.35 67.62 75.52
MoE-Adapters 37.38 93.85 80.51 68.31 81.89 84.07 90.01 73.98 90.55 67.67 66.40 75.87
CoLeCLIP 48.66 94.34 76.61 69.17 79.01 83.81 89.67 73.33 90.50 67.98 66.48 76.32
DPeCLIP 49.86 94.94 82.40 69.36 82.23 84.31 89.97 73.97 90.38 68.30 66.31 77.46
Table 2: $Last$, $Forgetting$, and $Avg$ results(%) in ODCL-TIL. CODA-Prompt and LAE are not applicable to the Avg metric.
Experiments
Experiments Setup
Datasets
The ODCL task comprises 11 datasets across various domains, including Aircraft (maji2013fine), Caltech101 (fei2004learning), CIFAR100 (2009Learning), DTD (cimpoi2014describing), EuroSAT (helber2019eurosat), Flowers (nilsback2008automated), Food (bossard2014food), MNIST (deng2012mnist), OxfordPet (parkhi2012cats), StanfordCars (krause20133d), and SUN397 (xiao2010sun). In the ODCL task, each dataset represents a distinct domain that needs to be addressed. The sequence in the ODCL task typically follows two orders, as previously adopted in the works of ZSCL (zheng2023preventing), ColeCLIP (li2024coleclip). The Order-I arranges the datasets alphabetically by their names, while the Order-II is determined randomly. The experimental results presented in the main text are based on the Order-I , whereas the supplementary materials provide results based on the Order-II.
Implementation Details
Building upon the previous research by ZSCL and ColeCLIP, we adopt the CLIP-ViT-B/16 (radford2021learning) model as the foundational architecture for tackling the ODCL task. All experimental settings strictly follow ColeCLIP (li2024coleclip), i.e., we employ the Adam optimizer with a fixed learning rate of $1\times 10^{-3}$, adopt a batch size of 128, and the specific number of epochs for each dataset can be found in the supplementary materials.
For the text encoder, we employ a single-layer learnable prompt as the general prompt integrated with a TSA module. The domain prior prompt generated by the TSA module replaces the original prompt in the $8$-th layer. For the image encoder, we utilize a 7-layer deep learnable prompts as general prompts, complemented by an ICA module. The instance domain prior prompt generated by the ICA module replaces the original prompt in the $8$-th layer. In line with the prompt length employed in ColeCLIP, we set the prompt length to 1 for both encoders. Similar to CLIP, the temperature $\tau$ is set to 0.01.
Evaluation Metrics
Following previous work such as ZSCL, ColeCLIP, we use Avg, Last, Transfer, and Forgetting as evaluation metrics to assess performance on the ODCL task. For the $n$th dataset $\mathcal{D}_{n}$, we use $A_{n}$, $F_{n}$ and $R_{n}$ to represent the mean accuracy, Frogetting and Transfer metrics, respectively.
Avg measures the overall classification accuracy across the 11 datasets, $Avg=\frac{1}{N}\sum_{n=1}^{N}A_{n}$, where $N$ is the total number of datasets. Specifically, after training on a given dataset, we evaluate the model’s performance on all 11 datasets (including both seen and unseen datasets) and compute the mean accuracy across these datasets, denoted as $A_{n}$, $A_{n}=\frac{1}{N}\sum_{u=1}^{N}A_{u}^{n}$, $A_{u}^{n}$ is the accuracy of the model on the $n$th dataset after training in $u$th dataset. After training on all 11 datasets, we calculate the mean of all accuracies to obtain Avg. Last refers to the final accuracy across the 11 datasets after training on the final task. Forgetting measures the degree of performance degradation on previously seen datasets. The computation involves calculating the mean accuracy for each previously seen dataset and then averaging these values to obtain the Forgetting metric, $Frogetting=\frac{1}{N}\sum_{n=1}^{N}F_{n}$ and $F_{n}=\frac{1}{N-n+1}\sum_{v=n}^{N}A_{v}^{n}$. Transfer assesses the model’s zero-shot capability on unseen datasets. It is computed by averaging the zero-shot accuracies of the datasets before the model is trained on them, $R_{n}=\frac{1}{n-1}\sum_{v=1}^{n-1}A_{v}^{n}$. The mean of these 10 values constitutes the Transfer metric, $Transfer=\frac{1}{N-1}\sum_{n=2}^{N}R_{n}$.
Method Caltech101 CIFAR100 DTD EuroSAT Flowers Food MNIST OxfordPet Cars SUN397 Mean
CLIP 88.19 67.31 44.68 55.26 70.22 88.50 59.45 89.04 64.71 65.16 69.25
ZSCL 84.56 67.45 44.77 51.54 69.02 87.60 62.34 87.08 59.65 66.41 68.04
MoE-Adapters 88.19 66.82 44.68 54.07 70.56 88.44 59.45 89.04 64.71 64.99 69.08
CoLeCLIP 88.19 65.12 44.68 54.10 68.82 88.49 59.45 89.04 64.71 65.11 68.77
DPeCLIP 88.19 67.18 44.68 53.97 70.58 88.19 59.45 89.04 64.71 64.78 69.08
Table 3: $Transfer$ results(%) in ODCL task.
[…]
- - - 61.00 57.87 59.01 69.25
$\checkmark$ 64.06 63.62 63.81 69.25
$\checkmark$ $\checkmark$ 75.20 83.41 83.79 69.07
$\checkmark$ $\checkmark$ 73.14 79.48 79.86 69.09
$\checkmark$ $\checkmark$ $\checkmark$ 76.05 84.64 85.05 69.08
Table 4: The impact of different module combinations on model performance. PTD is Prototype Task-ID discriminator, TC is text component and IC is image component.
[…]
$\checkmark$ 70.64 75.80 76.12 69.00
$\checkmark$ $\checkmark$ 71.76 77.22 77.75 68.90
$\checkmark$ $\checkmark$ 75.20 83.41 83.79 69.07
Table 5: Ablation analysis on the text branch. CoOp represents learnable prompt in $1$th layer. LP is learnable prompt in $8$th layer. TSA represents domain prior prompt in $8$th layer.
[…]
$\checkmark$ 70.15 74.84 75.16 68.12
$\checkmark$ $\checkmark$ 70.93 75.96 76.30 68.58
$\checkmark$ $\checkmark$ 73.14 79.48 79.86 69.09
Table 6: Ablation analysis on the image branch. VPT represents learnable prompt in 1-7 layers. LP is learnable prompt in $8$th layer. ICA represents instance-level domain prior prompt in $8$th layer.
Figure 3: (a) represents the influence of different prompt length and (b) represents the influence of different prompt replacement depth.
Comparison Methods
We selects CLIP (radford2021learning), ZSCL, MoE-Adapters (yu2024boosting), and ColeCLIP as the baseline methods for comparison. ZSCL represents the pioneering approach for ODCL task, MoE-Adapters explore the application of the MoE structure in the ODCL task while ColeCLIP builds upon it by further refining designs and enhancing performance. Since the MoE-Adapters does not fully align with this task, we modified its implementation to meet the demands of the ODCL task. The specific implementation details are provided in the supplementary materials. Following the methodology adopted by ColeCLIP, we also incorporate traditional continual learning methods, specifically CODA-Prompt (smith2023coda) and LAE (gao2023unified), for comparative analysis. However, since these methods are designed for closed-domain scenarios and lack zero-shot capabilities, they are only applicable for evaluating the Last and Forgetting metrics, and not for assessing the Avg and Transfer metrics.
Main Properties
Performance on ODCL-CIL
As previously mentioned, ODCL-CIL presents a more challenging task compared to ODCL-TIL due to the absence of Task-ID information during the inference stage. Table 1 provides a comparison between our proposed method DPeCLIP and other approaches, highlighting significant improvements in performance.
Notably, DPeCLIP achieves improvements of 4.90% and 3.33% over CoLeCLIP in the Last and Forgetting metrics, respectively. For the Caltech101 and CIFAR100 datasets, which are significantly affected by the category-relatedness issue, the first category of methods, ZSCL and ColeCLIP, exhibited noticeable forgetting phenomena. MoE-Adapters and our method demonstrate that a robust Task-ID discriminator can effectively address the category-relatedness problem. Additionally, due to the shortcomings of MoE-Adapters’ Task-ID discriminator, the performance on many datasets is significantly affected, such as Caltech101, EuroSAT, and Food.
Our method shows an enhancement of 2.37% compared to CoLeCLIP in Avg. Although MoE-Adapters demonstrates strong performance on the DTD task, it utilizes the AutoEncoder as the Task-ID discriminator, which fails to recognize test images from EuroSAT. Thus, it negatively impacts the overall performance. In comparison, our prototype Task-ID discriminator approach demonstrates exceptional performance in terms of Task-ID identification accuracy, which ultimately enhances the overall model performance. Overall, our method demonstrates robust performance, consistently outperforming previous approaches across the majority of datasets.
[…]
Domain-level 67.11 62.32 69.18
Category-level 76.05 84.65 85.05
Table 7: The impact of prototype granularity for the Task-ID discriminator. Domain-level granularity represents one prototype for each domain, while category-level granularity represents one prototype for each category within each domain.
[…]
TP 69.39 72.27 73.91
IP 76.00 84.39 85.03
TP&IP 76.05 84.65 85.05
Table 8: The impact of different types of prototypes on the Task-ID discriminator. TP represents text prototypes, while IP represents image prototypes.
[…]
IP 69.05 75.92 84.47 84.89 77.35 86.72 86.72
TP 69.01 75.95 84.55 84.95 77.36 86.77 86.77
IP&TP 69.08 76.05 84.65 85.05 77.46 86.89 86.89
Table 9: The impact of different type of prototype for domian prior prompt. TP is referring text prototype and IP is referring image prototype.
Figure 4: Comparison of t-SNE between MoE-Adapters and our method for the Task-ID discriminator’s output.
Performance on ODCL-TIL
For the ODCL-TIL task, DPeCLIP continues to demonstrate exceptional performance, as illustrated in Table 2. In terms of the Last, Forgetting, and Avg metrics, DPeCLIP achieves the highest results, surpassing CoLeCLIP by 2.18%, 1.74%, and 1.14%, respectively. This highlights that DPeCLIP not only excels in addressing the issue of forgetting but also achieves superior learning outcomes across various domain tasks. DPeCLIP effectively balances the trade-off between learning capability and forgetting mitigation, resulting in outstanding overall performance.
Transfer performance
Table 3 presents the Transfer metric for models on the ODCL-TIL task. ZSCL, by fully fine-tuning the CLIP model, experiences a significant decline in transfer performance, with a reduction of 1.21%. CoLeCLIP mitigates this issue by maintaining a large vocabulary, resulting in notable improvements. MoE-Adapters utilizes a mixture of expert structure to retain CLIP’s original knowledge, and it exhibits relatively good performance on the transfer metric. In contrast, DPeCLIP introduces domain prototypes as input for the domain prior prompts, aiming to retain the original knowledge of CLIP as much as possible, thereby concurrently outperforming other methods on various other metrics, achieving the same level of performance on the Transfer metric as MoE-Adapters, with only a 0.17% reduction.
Analysis and discussion
Module Analysis
We conduct comprehensive ablation experiments on DPeCLIP to demonstrate the impact of different components of our proposed DPeCLIP. Since the ODCL-CIL task can better evaluate the overall performance of the model, Table 4 shows the performance using different components on ODCL-CIL task. The first row represents the performance of the original CLIP on ODCL-CIL task. By incorporating the prototype Task-ID discriminator, without requiring any training, the Avg, Last, and Forgetting metrics show significant improvements, 3.06%, 5.75% and 4.8% respectively. This method effectively addresses the issue of category-relatedness across different domains.
TC (Text Component) refers to the general prompt and domain prior prompt generated by the TSA module for the text encoder and IC (Image Component) represents the general prompt and instance-level domain prior prompt generated by the ICA module for the image encoder. Compared to IC, the TC shows a more significant increase in the three main metrics, demonstrating that fine-tuning the text component of CLIP is more effective. Additionally, the domain prior prompt generated from the domain prototype performs well in retaining the original knowledge of CLIP, resulting in only a slight decline in the Transfer metric. The integration of the three modules achieves the best performance in Avg, Last, and Forgetting metrics, with only a very slight decline in the Transfer metric. The detailed ablation experiments for each component are shown below.
Detailed ablation study of TC
Table  5 shows a detailed analysis of TC. CoOp represents the general prompt for the first layer, LP refers to the learnable prompt for the $8$th layer, and TSA module provides the domain prior prompt for the $8$th layer. Compared to the combination of CoOp and LP, the combination of CoOp and TSA shows significant improvements across all metrics, demonstrating the effectiveness of the domain prior prompt for the text encoder.
Detailed ablation study of IC
Table 6 presents a detailed analysis of IC. VPT represents the learnable prompt for the first $7$ layers used as the general prompt, LP refers to the learnable prompt for the $8$th layer, and ICA stands for the instance-level domain prior prompt generated by the ICA module for the $8$th layer. Notably, the combination of VPT and ICA shows significant growth not only in all metrics, including the transfer metric, demonstrating that the domain prior prompt retains more of CLIP’s original knowledge compared to the learnable prompt.
Prompt Length
Figure 3(a) presents the performance of DPeCLIP with different prompt length on ODCL task. Given the same settings for hyperparameters such as training iterations and learning rate, DPeCLIP exhibits the best performance across all metrics when the prompt length is 1.
Replacement depth
Figure 3(b) presents the replace depths of the domain prior prompt obtained using the TSA and ICA modules. We conducted experiments by using layers 5 to 11 of the CLIP encoder as the replacement depth. Although the Transfer metric achieves the best performance when inserted at layer 6, a comprehensive evaluation leads us to ultimately select layer 8, as it yields the best overall performance. We hypothesize that inserting the layer 8 can enable the model to better balance the consideration of low-level and high-level features.
Prototype granularity for the Task-ID discriminator
Table 7 shows the impact of different prototype granularities on the Task-ID discriminator. Compared to domain-level prototypes, category-level prototypes can retain more domain information, resulting in more accurate Task-ID judgments and better model performance.
Prototype type for the Task-ID discriminator
Table 8 shows the impact of different types of prototypes on the Task-ID discriminator. Compared to the confusion issues associated with TP, IP better represents domain-level information, resulting in improved performance of the Task-ID discriminator. Additionally, we found that using IP and TP together can slightly enhance overall performance.
Prototype type for the domain prior prompt
Table 9 illustrates the influence of different type of prototype for domain prior prompt. It can be observed that using either type of prototype individually can achieve good performance. However, since combining both yields the best results, we ultimately choose to use a combination of TP and IP.
t-SNE Visualization
Figure 4 shows the different distributions between MoE-Adapters and our method in the Task-ID discriminator’s output. MoE-Adapters exhibit significant confusion among different datasets, whereas our method effectively reduces this confusion. The comparison of the magnified sections further demonstrates the superior performance of our method on the Task-ID discriminator task.
Conclusion
In this work, we propose DPeCLIP to address the challenges faced by VLMs in ODCL task. By proposing the methods of Domain Prior Classifier and Domain Prior Prompt, our model achieves state-of-the-art performance in both ODCL-CIL and ODCL-TIL settings. However, there remains room for improvement in the model’s accuracy in identifying dataset Task-ID and in preserving the original zero-shot capabilities of the model. Future work will focus on addressing these issues.